# HAF-RM: A Hybrid Alignment Framework for Reward Model Training

### v2

@@ Abstract
The reward model has become increasingly important in alignment, assessment, and data construction for large language models (LLMs). Most existing researchers focus on enhancing reward models through data improvements, following the conventional training framework for reward models that directly optimizes the predicted rewards. In this paper, we propose a hybrid alignment framework HaF-RM for reward model training by introducing an additional constraint on token-level policy probabilities in addition to the reward score. It can simultaneously supervise the internal preference model at the token level and optimize the mapping layer of the reward model at the sequence level. Theoretical justifications and experiment results on five datasets show the validity and effectiveness of our proposed hybrid framework for training a high-quality reward model. By decoupling the reward modeling procedure and incorporating hybrid supervision, our HaF-RM framework offers a principled and effective approach to enhancing the performance and alignment of reward models, a critical component in the responsible development of powerful language models. We release our code at https://haf-rm.github.io.
@@ 1 Introduction
Recent periods have witnessed a continuous evolution of Large Language Model (LLM) techniques, especially pre-training Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020) and instruction tuning Wei et al. (2021); Wang et al. (2022); Yue et al. (2023). Researchers start to shift their focus from generating correct responses to aligning responses more closely with human preferences (Russell, 2014). As an efficient alternative to human feedback, the reward model for generative language models emerges, facilitating scalable alignment in training (Christiano et al., 2017; Stiennon et al., 2020), response generation (Gao et al., 2023; Mudgal et al., 2024; Jinnai et al., 2024), data construction(Yuan et al., 2023) etc.
Despite the availability of numerous sophisticated reward models (Kopf et al., 2023; Zhu et al., 2023), these exist several key limitations. First, most reward models originate from industry and are not open-source, making further training and transfer impossible. Second, prior studies have highlighted incorrect and ambiguous preferences within the training data of these reward models (Bai et al., 2022; Pitis, 2023). These two issues both limit the quality and generalizability of existing reward models, necessitating further enhancement either from the data perspective or the training process. While recent researches mainly focus on enriching data sources for better reward models, including utilizing external tools or information sources to enhance generalization (Li et al., 2023a; Sun et al., 2023) or leveraging fine-grained signals (Wu et al., 2023; Cao et al., 2024) and their combinations (Go et al., 2023; Lai et al., 2024), we focus on the training framework of reward models in this work.
@@
A reward model is typically structured with two components: a transformer-based model (referred to as the internal preference model) that outputs preference vectors for each token, and a projection module called “reward layer” (usually a linear layer with normalization) that maps these vectors to sequence-level rewards. The standard practice for training the reward model involves utilizing the ranking loss of paired rewards. However, optimizing both two components using such a single sequence-level objective may cause insufficient supervision for token-level preference modeling. We argue that hybrid optimization of the two components of the reward model with corresponding token-level and sequence-level objectives will lead to more consistent improvement.
Since a policy model is also based on an internal preference model to predict the expected reward for each action/token, essentially acting as a Q-function under token-level supervision Rafailov et al. (2024), we propose a Hybrid Alignment Framework (HAF). This framework jointly optimizes the reward model and policy model with a shared internal preference model. With the policy loss, we can directly supervise the internal preference model at the token level while simultaneously optimizing the mapping layer of the reward model using the reward loss, enabling more effective alignment of the reward model.
We provide both theoretical justifications and empirical experiments to demonstrate the effectiveness of our HAF. In the experiment section, we compare the performance of reward models trained using our framework against those resulting from baseline approaches across four public datasets. The results highlight the advantage of HAF with different policy losses integrated. Further analysis reveals that using additional policy loss can improve the performance of policy model calibration, which opens a new horizon for training high-quality reward models.
2 Preliminary
@@ 2.1 Notation
  • The oracle (optimal) value is denoted as the corresponding letter with an asterisk such as $\bm{r}^{*}$(oracle reward model), $\phi^{*}$(optimal model preference), $\mathrm{F}^{*}$(optimal reward prediction layer) and $\mathrm{K}^{*}$(optimal policy prediction layer).
2.2 Training Loss
@@
$\mathcal{L}_{s}\coloneqq\underset{d\sim\mathcal{P}}{\mathbb{E}}\left[{\mathrm{D}_{1}\left({\bm{r}\left({d}\right),\bm{r}^{*}\left({d}\right)}\right)}\right]$ (1)
For notational convenience, we use $d$ to denote $\left({x,y}\right)$ and use $\underset{\bm{r}}{\mathrm{argmin}}\mathcal{L}_{s}$ or $\underset{\mathrm{F},\phi}{\mathrm{argmin}}\mathcal{L}_{s}$ to represent the model training with the standard reward loss.
@@
Similar to the reward loss, standard policy loss aims to measure the error of the policy model.
$\mathcal{L}_{P}\coloneqq\underset{d\sim\mathcal{P}}{\mathbb{E}}\left[{\mathrm{D}_{2}\left({\bm{\pi}\left({d}\right),\bm{\pi}^{*}\left({d}\right)}\right)}\right]$ (2)
@@
$\mathcal{L}_{H}\coloneqq\underset{d\sim\mathcal{P}}{\mathbb{E}}\left[\mathrm{D}_{1}\left({\bm{r}\left({d}\right),\bm{r}^{*}\left({d}\right)}\right)\right.$ (3)
@@
$=\underset{d\sim\mathcal{P}}{\mathbb{E}}\left[\mathrm{D}_{1}\left({\mathrm{F\circ\phi}\left({d}\right),\mathrm{F^{*}\circ\phi^{*}}\left({d}\right)}\right)\right.$
@@
where $\alpha$ is a hyperparameter to balance losses from the reward and policy model, $\phi$ is the shared internal preference model which receives gradients from both loss terms. Similarly, $\underset{\mathrm{F,K},\phi}{\mathrm{argmin}}\mathcal{L}_{H}$ and $\underset{\bm{r,\pi}}{\mathrm{argmin}}\mathcal{L}_{H}$ represent the model training with our hybrid alignment loss.
3 Hybrid Alignment Framework
3.1 Model Implementation
@@
The most commonly used decoder-only LLM consists of stacked transformer blocks (Vaswani et al., 2017) or similar structures, and a linear layer for policy projection. In the reward model, only the shape of the final linear layer is adjusted to match the format of the reward value output compared to the policy model. We retain two linear layers for our model, enabling it to output rewards and probabilities simultaneously.
To significantly reduce the resources required for training, it is standard practice to initialize the internal preference module of the reward model with a fine-tuned language model as it retains the model’s language modeling capabilities.
3.2 Loss Calculation
There is consensus on the specific calculation method for the reward loss. In avoiding the issue of uncertain reward values, the Bradley-Terry model (Christiano et al., 2017) is used to transform the reward modeling problem into a probability optimization problem. Treating the problem as a binary classification task yields the popular form of reward loss function:
$\mathcal{L}_{s}\;=\,\quad\underset{d\sim\mathcal{P}}{\mathbb{E}}$ $\left[\mathrm{D}_{1}\left({\bm{r}\left({d}\right),\bm{r}^{*}\left({d}\right)}\right)\right]$ (4)
$=\underset{\left({x,y,y^{\prime}}\right)\sim\mathcal{D}}{\mathbb{E}}$ $\left[-\log\sigma\left({\bm{r}\left({x,y}\right)-\bm{r}\left({x,y^{\prime}}\right)}\right)\right]$
where $\sigma\left({\cdot}\right)$ is the sigmoid function.
Given the preference data, there currently does not exist a universally optimal policy loss. However, since the derivation of the DPO loss is based on assumptions similar to those made for the reward loss (as detailed in Appendix C.2), we choose to use the DPO loss as the method for calculating the policy loss.
$\mathcal{L}_{P}\;=\,\quad\underset{d\sim\mathcal{P}}{\mathbb{E}}$ $\left[{\mathrm{D}_{2}\left({\bm{\pi}\left({d}\right),\bm{\pi}^{*}\left({d}\right)}\right)}\right]$ (5)
$=\underset{\left({x,y,y^{\prime}}\right)\sim\mathcal{D}}{\mathbb{E}}$ $\left[{-\log\sigma\left({\tau\left({pd_{win}-pd_{lose}}\right)}\right)}\right]$
where
Combining the two losses, we have our HaF loss calculate in the following manner:
$\mathcal{L}_{H}=\mathcal{L}_{s}+\alpha\cdot\mathcal{L}_{P}$ (6)
We will elaborate in Appendix C.1 on why Eq. 4 and Eq. 5 hold and why there is no optimal model on the right-hand side.
3.3 Theoretical Analysis
In this subsection, we present several properties of HaF that are independent of the specific calculation methods of the two loss functions. We will start from Section 2.2.
In practice, functions such as $\mathrm{F}$ and $\phi$ are represented by parameterized models with finite parameters, and thus cannot precisely model arbitrary distributions. Here we show that under certain assumptions, using the hybrid alignment loss can yield a better solution than simply using the standard reward loss.
Proposition 1.
Unless $K$ can exactly fit $K^{*}$, there exists $\epsilon>0$, such that
holds for all $\alpha\in(0.1,2)$, where $K_{H},\phi_{H}=\underset{\mathrm{K},\phi}{\mathrm{argmin}}\mathcal{L}_{H}$ in Equation 3 and $\phi_{s}=\underset{\phi}{\mathrm{argmin}}\mathcal{L}_{s}$ in Equation 4.
Here we use $\mathrm{argmin}$ to represent the best models optimized with the corresponding loss functions, so $\phi_{H}$ and $\phi_{s}$ are not equal to $\phi^{*}$ although $\phi^{*}$ is the minimum mathematically. Intuitively this indicates that the model learned from the joint calibrated loss outperforms the one learned solely from the preference space using the standard reward loss.
Proposition 2.
The detailed derivations for both propositions are provided in Appendix D. Here we obtain an upper bound on the model preference error. By tuning the hyperparameter $\alpha$, the right term can be strictly negative. In other words, model preference space trained with our calibrated loss can be strictly closer to the true preference space compared to the standard reward loss. (In practice, there is no need for an exhaustive search, we find $\alpha=0.2$ already yields satisfactory results. We give a discussion about this in B)
4 Experiment setup
@@ 4.1 Datasets
We comprehensively assess the performance of our framework using five public datasets, namely Anthropic-HH-Harmless (HH-harmless) (Bai et al., 2022), Anthropic-HH-Helpful (HH-Helpful) (Bai et al., 2022), Beaver Safe (BS) (Ji et al., 2023), Alpaca Human Pref (AHP) (Dubois et al., 2023) and Chatbot Arena (CA) (Zheng et al., 2023). Note that AHP and CA do not have original data split for evaluation, we randomly extract 10% from the original data as a test set, the details of the used datasets are shown in Tab 1.
Name Size Words/QA Tokens/QA
@@
Table 1: Statistics of the Training Datasets
4.2 Comparative Models
@@
We compare our framework with the standard training approach, in which the reward model only has a reward layer for reward prediction and is optimized via Eq. 4.
@@
Although DPO loss (Eq. 5) is typically used for training policy models rather than reward models, it can implicitly convert the model’s outputs into reward values (Rafailov et al., 2023). Therefore, the DPO model can also be considered a reward model (Rafailov et al., 2024). Following the work of Lambert et al. (2024), we also evaluate the model trained with DPO loss.
@@
Under our framework, the reward model has both the reward and policy layer for predicting sequence-level rewards and providing token-level probabilities.
In our implementation, we use Phi-2-2.7B and Mistral-7B-Instruct-v0.2 as our base model. We train Phi-2 and Mistral-7B using full-parameter and Low-rank Adaptation (LoRA) Hu et al. (2022), respectively. More experiment setup can be found in Appendix A.
@@ 5.1 Intrinsic performance of Reward Models
The primary function of a reward model is to evaluate the quality of responses to a given question, which involves accurately comparing two answers to the same question. Using judgment accuracy as the evaluation metric, we conduct several experiments to assess the effectiveness of HaF in training the reward model.

### v3

@@ Abstract
The reward model has become increasingly important in alignment, assessment, and data construction for large language models (LLMs). Most existing researchers focus on enhancing reward models through data improvements, following the conventional training framework for reward models that directly optimizes the predicted rewards. In this paper, we propose a hybrid alignment framework HaF-RM for reward model training by introducing an additional constraint on token-level policy probabilities in addition to the reward score. It can simultaneously supervise the internal preference model at the token level and optimize the mapping layer of the reward model at the sequence level. Experiment results on five datasets sufficiently show the validity and effectiveness of our proposed hybrid framework for training a high-quality reward model. By decoupling the reward modeling procedure and incorporating hybrid supervision, our HaF-RM framework offers a principled and effective approach to enhancing the performance and alignment of reward models, a critical component in the responsible development of powerful language models. We release our code at https://haf-rm-anonymized.github.io.
@@ 1 Introduction
Recent periods have witnessed a continuous evolution of Large Language Model (LLM) techniques, especially in pre-training Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020) and instruction tuning Wei et al. (2021); Wang et al. (2022); Yue et al. (2023). As these models advance, researchers have shifted their focus from generating correct responses to aligning outputs more closely with human preferences (Russell, 2014) through Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022). As an efficient alternative to human feedback, reward models for generative language models emerge, facilitating scalable alignment in training (Christiano et al., 2017; Stiennon et al., 2020), response generation (Gao et al., 2023; Mudgal et al., 2024; Jinnai et al., 2024), and data construction(Yuan et al., 2023) etc.
Despite the availability of numerous sophisticated reward models (Kopf et al., 2023; Zhu et al., 2023), several key limitations remain. First, many reward models are proprietary and closed-source, originating from industry, which restricts their further training and transfer. Second, prior studies have highlighted incorrect and ambiguous preferences within the training data of these reward models (Bai et al., 2022; Pitis, 2023). These two issues both limit the quality and generalizability of existing reward models, necessitating further enhancement either from the data perspective or the training process. Recent efforts primarily focus on enriching data sources to improve reward models, including incorporating external tools or information sources to enhance generalization (Li et al., 2023a; Sun et al., 2023) or leveraging fine-grained signals (Wu et al., 2023; Cao et al., 2024) and their combinations (Go et al., 2023; Lai et al., 2024). In contrast, this work aims to improve the training framework of reward models.
@@
A reward model is typically structured with two components: a transformer-based model (referred to as the internal preference model), and a projection module called “reward layer” (usually a linear layer with normalization). The former outputs preference vectors for each token, while the latter maps these vectors to sequence-level rewards. The standard practice for training the reward model involves utilizing the ranking loss of paired rewards. However, optimizing both two components using such a single sequence-level objective may cause insufficient supervision for token-level preference modeling. We argue that hybrid optimization of the two components of the reward model with corresponding token-level and sequence-level objectives will lead to more consistent improvements.
Given that a policy model also relies on an internal preference model to predict expected rewards for each action or token, essentially acting as a Q-function under token-level supervision Rafailov et al. (2024), we propose a Hybrid Alignment Framework (HAF). This framework jointly optimizes the reward model and the policy model by sharing the internal preference model. With an additional policy loss, we can directly supervise the internal preference model at the token level, while simultaneously optimizing the mapping layer of the reward model using the reward loss, enabling more effective alignment of the reward model.
We provide massive empirical experiments with an intuitional justification to demonstrate the effectiveness of our HAF. In the experiment section, we compare the performance of reward models trained using our framework against those resulting from traditional baseline and DPO approaches across five public datasets. The results highlight the advantage of HAF with different policy losses integrated. Further analysis reveals that using additional policy loss can improve the performance of policy model calibration, which opens a new horizon for training high-quality reward models.
2 Hybrid Alignment Framework
In this section, we first introduce the necessary notations (Section 2.1). Then we derive the formation of reward loss and policy loss as well as their practical calculation methods (Section 2.2), and propose HaF to effectively utilize the similarity between the reward model and the policy model (Section 2.3). Finally, we provide an intuition-based explanation for why HaF works (Section2.4).
2.1 Notation
@@
  • The Oracle value is denoted as the corresponding letter with an asterisk such as $\bm{r}^{*}$(Oracle reward model), $\phi^{*}$(Oracle model preference), $\mathrm{F}^{*}$(Oracle reward prediction layer) and $\mathrm{K}^{*}$(Oracle policy prediction layer).
2.2 Basic Loss Functions
@@
[…]
We use $d$ to denote $\left({x,y}\right)$ for notational simplicity.
In avoiding the issue of uncertain reward values, there is consensus on the use of the Bradley-Terry model (Bradley and Terry, 1952) to transform the reward modeling problem into a probability optimization problem (Stiennon et al., 2020; Rafailov et al., 2023; Meng et al., 2024), which yields the popular form of a binary classification cross-entropy loss:
$\mathcal{L}_{s}\leftarrow\underset{\left({x,y,y^{\prime}}\right)\sim\mathcal{D}}{\mathbb{E}}$ $\left[-\log\sigma\left({\bm{r}\left({x,y}\right)-\bm{r}\left({x,y^{\prime}}\right)}\right)\right]$ (2)
where $\sigma\left({\cdot}\right)$ is the sigmoid function (derivation can be found in C.1).
@@
Similar to the reward loss, the standard policy loss aims to measure the error of the policy model.
[…]
Here, we use DPO Rafailov et al. (2023) for calculating policy loss since its derivation is similar to that made for the reward loss (as detailed in Appendix C.2).
$\mathcal{L}_{P}\leftarrow\underset{\left({x,y,y^{\prime}}\right)\sim\mathcal{D}}{\mathbb{E}}$ $\left[{-\log\sigma\left({\tau\left({pd_{win}-pd_{lose}}\right)}\right)}\right]$ (4)
Figure 2: HaF tends to assign higher scores to the responses it generates. The x-axis represents the score difference between the ideal reward model’s evaluation of the content generated by HaF’s policy head and the content generated by the model trained with DPO. The y-axis indicates the score difference when HaF evaluates these two outputs. Different colors represent different model checkpoint selection strategies.
2.3 HaF Implementation
@@
[…]
@@
[…]
@@
where $\alpha$ is a hyperparameter to balance losses from the reward and policy model, $\phi$ is the shared internal preference model which receives gradients from both loss terms.
Model structure
@@
The most commonly used decoder-only LLM consists of stacked transformer blocks (Vaswani et al., 2017) or similar structures, and a linear layer for policy projection. In the reward model, only the shape of the final linear layer is adjusted to match the format of the reward value output compared to the policy model (Stiennon et al., 2020). We retain two linear layers for our model, enabling it to output rewards and probabilities simultaneously, as shown in Figure 1.
2.4 Why HaF is Better?
Figure 2 shows the consistency between the reward model and the policy model in preference learning. Despite possessing similar generation quality, the policy model which shares parameters with the reward model is rated higher, indicating that the two models do have resembling preferences when they have the same internal preference model. We will elaborate on this finding in Appendix E.1.
Besides, we provide an intuitive explanation of why the hybrid alignment loss can yield a better solution than simply using the standard reward loss.
Claim 1.
The model learned from the joint calibrated loss outperforms the one learned solely from the preference space using the standard reward loss. Details can be found in Appendix D.
Claim 2.
Policy loss can act as a regularization term preventing the inner representation from degrading, so HaF tends to outperform the traditional training framework.
3 Experimental Setup
@@ 3.1 Datasets
We comprehensively evaluate the performance of our framework using five public datasets: Anthropic-HH-Harmless (HH-harmless) (Bai et al., 2022), Anthropic-HH-Helpful (HH-Helpful) (Bai et al., 2022), Beaver Safe (BS) (Ji et al., 2023), Alpaca Human Pref (AHP) (Dubois et al., 2023), and Chatbot Arena (CA) (Zheng et al., 2023). Since AHP and CA do not provide original data split for evaluation, we randomly extract 10% from the original data as the test set. Detailed statistics of our used datasets for training are shown in Table 1.
Dataset Size #Word/QA #Token/QA
@@
Table 1: Statistics of the Training Subsets.
3.2 Compared Models
@@
First, we compare our framework with the standard training approach, wherein the reward model only has a reward layer dedicated to reward prediction and is optimized only with reward loss, as delineated in Eq. 2.
@@
Although DPO loss (as in Eq. 4) is typically used for training policy models rather than reward models, it can implicitly convert model’s outputs into reward values (Rafailov et al., 2023). Therefore, the DPO-trained model can also function as a reward model (Rafailov et al., 2024). Following the work of Lambert et al. (2024), we also evaluate the model trained with DPO loss.
@@
Under our framework, the reward model has both a reward layer and a policy layer for predicting sequence-level rewards and providing token-level probabilities.
Our framework is implemented based on three different backbone LLMs including both pre-trained and fine-tuned models: Phi-2-2.7B Javaheripi et al. (2023), Mistral-7B-base-v0.3 and Mistral-7B-Instruct-v0.2 Jiang et al. (2023). We train Phi-2 and Mistrals using full-parameter and Low-rank Adaptation (LoRA) Hu et al. (2022) strategies, respectively. More implementation details can be found in Appendix A.
@@ 4.1 Intrinsic Performance of Reward Models
The primary function of a reward model is to evaluate the quality of responses to a given question, which involves accurately comparing pairs of answers to the same question. To demonstrate the effectiveness of our HaF in training reward models, we first conduct several experiments evaluating the intrinsic performance of our trained reward model, specifically by taking judgment accuracy as the evaluation metric.
@@ 4.1.1 Overall Performance
DPO(Mistral-base) 64.60 69.90 68.80 91.70 53.80 69.76
Baseline(Mistral-base) 72.60 69.80 64.20 78.30 50.40 67.06
HaF (Mistral-base) 73.00 70.00 74.40 85.40 56.30 71.82